# Bridged Clustering for Representation Learning: Semi-Supervised Sparse Bridging

### v1

@@ 8 Conclusion & Future Work
Bridged Clustering offers a new and simple vantage for SSL by elevating output-only data from a peripheral role to a central asset. By decoupling input from output representation and stitching them together with a sparse, interpretable bridge, our framework addresses a gap between classic input-centric SSL and the realities of modern, richly available but poorly aligned data ecosystems.
Our analysis shows that Bridged Clustering can become a powerful predictor with bounded mis-clustering and mis-bridging rates. Empirically, it performs competitively among strong baselines across diverse domains and data modalities. The method is model-agnostic, computationally efficient, and supports bidirectional inference without retraining. These properties suggest a broader research direction: learning from abundant unpaired observations by first modeling each view independently, and then bridging them with minimal supervision.
Future work can explore domains with overlapping or hierarchical categories, extending to inductive settings, incorporating soft or probabilistic alignments, and integrating with unsupervised techniques such as manifold learning or self-supervised embeddings. These lead to more flexible SSL systems that can effectively harness the rich but misaligned data landscapes common in modern applications.
9 Ethics statement
@@
10 Reproducibility statement
We have taken several steps to ensure reproducibility. All datasets used in our experiments are publicly available, and preprocessing scripts are included in our supplementary materials. The implementation of Bridged Clustering, including clustering, alignment, and evaluation routines, is fully documented and released alongside the paper.
Hyperparameters, random seeds, and experimental configurations for all reported results are explicitly listed in the code and appendix. Each experiment is repeated across multiple random seeds to quantify variability, and results are reported with appropriate statistics.
Additionally, we provide detailed instructions for reproducing both the main figures and tables, including cluster visualizations, MSE distributions, and win-rate calculations. This enables other researchers to verify our findings, test alternative design choices, and extend Bridged Clustering to new datasets or tasks with minimal ambiguity.
@@ A.1 BIOSCAN-5M
To convert the raw data into cluster-able data that we can easily train and evaluate on, we encode the specimens images with a EfficientNetB0 image encoder (Tan & Le, 2019) and the DNA barcodes with BarcodeBERT (Arias et al., 2023). We also narrow our sampling to the Insecta Class since it contains 98% of the BIOSCAN-5M data.
@@ A.2 WIT
To curate our dataset, we select the first 100,000 samples from WIT, encoding the images with ResNet50 (He et al., 2016)and the text captions with Sentence-BERT(Reimers & Gurevych, 2019).
@@ Appendix C Bridged Clustering Algorithm (Full version with Inverse Prediction)
sparse supervised set $\mathcal{S}=\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{k}$; Input Test-set $\mathcal{X_{\text{test}}}=\{x_{1},x_{2},\ldots,x_{|\mathcal{X_{\text{test}}}|}\}$; Output Test-set $\mathcal{Y_{\text{test}}}=\{y_{1},y_{2},\ldots,y_{|\mathcal{Y_{\text{test}}}|}\}$ (Note: $\mathcal{X_{\text{test}}}=\mathcal{X}$, $\mathcal{Y_{\text{test}}}=\mathcal{Y}$ in transductive setting)
@@
Appendix D Detailed Experiment Results: MSE
@@
Appendix E Detailed Experiment Results: Cluster Quality and Accuracy (Transductive)
@@
Appendix F LLM Usage
@@
Appendix G More Information

### v3

@@ 8 Conclusion & Future Work
Bridged Clustering offers a simple new perspective for SSL by elevating output-only data from a peripheral role to a central asset. By decoupling input and output representations and connecting them via a sparse, interpretable bridge, our framework addresses a gap between classic input-centric SSL and modern data ecosystems with abundant but misaligned observations.
Empirically, Bridged Clustering performs competitively across diverse domains and modalities. The method is model-agnostic, computationally efficient, and supports bidirectional inference without retraining. By separating representation learning from supervision, it uncovers latent structure in both input and output spaces and aligns them with minimal labels. Concept-level alignment between independently clustered spaces provides a strong inductive bias, enabling generalization without dense supervision, which is especially valuable in settings with limited labeled data, high domain variability, or distribution shifts.
Our framework is naturally extensible while preserving interpretability and efficiency. Extending our model to algorithmic variants such as Soft-Bridging and Cluster-wise Regression, brings robustness in adversarial data conditions. Future work can explore combination of Bridged Clustering with unsupervised techniques like manifold learning or self-supervised embeddings, in effort to build more flexible semi-supervised learning systems capable of leveraging richly available but misaligned data.
Our findings suggest a future where modular, concept-aligned learning can extract meaningful structure from unsupervised but richly available observations, without the need for dense labels or heavy models.
Impact Statement
@@ A.1 BIOSCAN-5M
To convert the raw data into cluster-able data that we can easily train and evaluate on, we encode the specimens images with a EfficientNetB0 image encoder (Tan and Le, 2019) and the DNA barcodes with BarcodeBERT (Arias et al., 2023). We also narrow our sampling to the Insecta Class since it contains 98% of the BIOSCAN-5M data.
@@ A.2 WIT
To curate our dataset, we select the first 100,000 samples from WIT, encoding the images with ResNet50 (He et al., 2016)and the text captions with Sentence-BERT(Reimers and Gurevych, 2019).
@@ Appendix C Bridged Clustering Algorithm (Full version with Inverse Prediction)
sparse supervised set $\mathcal{S}=\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{k}$; Input Test-set $\mathcal{X_{\text{test}}}=\{x_{1},x_{2},\ldots,x_{|\mathcal{X_{\text{test}}}|}\}$; Output Test-set $\mathcal{Y_{\text{test}}}=\{y_{1},y_{2},\ldots,y_{|\mathcal{Y_{\text{test}}}|}\}$ (Note: $\mathcal{X_{\text{test}}}=\mathcal{X}$, $\mathcal{Y_{\text{test}}}=\mathcal{Y}$ in the transductive setting)
@@
Appendix D Runtime and Memory Complexity
$\text{training}=f(n_{X},C,d)+f(n_{Y},C,d^{\prime})+O(k),$
$\text{per-query inference}=O(Cd+d^{\prime}),$
Method Training time Memory
EOT $O(S\,n_{X}n_{Y})$ $O(n_{X}n_{Y})$
GW $O(T_{\text{GW}}(n_{X}^{2}n_{Y}+n_{X}n_{Y}^{2}))$ $O(n_{X}^{2}+n_{Y}^{2}+n_{X}n_{Y})$
Figure 3: Empirical Runtime shows linear growth, mirroring theoretical bound (Inductive Experiments).
Appendix E Detailed Experiment Results: MSE
@@
Appendix F Detailed Experiment Results: Cluster Quality and Accuracy (Transductive)
@@
Appendix G Experiment Results for Algorithmic Variants
Figure 14: Inductive Experiments using Algorithmic Variants: The 4 distribution plots of the same color represent the settings with 1,2,3,4 supervised samples per cluster. Notice that Soft-Bridging is the overall strongest method in addressing both overlapping and misccounted clusters. For Cluster-wise Regression, Supervised Refinement alone performs unstably, but adding centroid to regressor fitting improves model performance, slightly outperforming the original Bridged Clustering model. Here, miscounting refers to overcounting the number of clusters, and we note that undercounting experiments show similar results.
Figure 15: Transductive Experiments using Algorithmic Variants. Similar Observations with Inductive Setting.
Appendix H Experiment Results for Adversarial Conditions
Our base model Bridged Clustering does incur more loss in cluster-adversarial conditions, but still maintains overall competitive performance, among the strongest models with low MSE.
Figure 16: Running main experiment on COCO with Overlapping Data Classes
Figure 17: Running main experiment on Large Cluster Counts for BIOSCAN
Figure 18: Running main experiment on Misscounted Cluster numbers for COCO and Reversed COCO (Shown inductive results, but transductive outcomes are similar. Miscounting as shown here refers to overcounting the number of clusters, and we note that undercounting experiments also show similar results.)
Appendix I Using Different Encoders and Clustering Methods
Figure 19: Using Different Image Encoders for BIOSCAN Experiment
Figure 20: Using Different Clustering Methods for COCO Experiment
Appendix J Illustrations for Bridged Clustering
Figure 21: Illustrations for Bridged Clustering in BIOSCAN. In both Image and DNA spaces, datapoints from same species are often clustered together due to embedding proximity. Highlighted datapoints are supervised.
Figure 22: Illustrations for Bridged Clustering in COCO. In both Image and text spaces, datapoints from similar categories are often clustered together due to embedding proximity. Highlighted datapoints are supervised.
Appendix K BLEU Score for Image-Captioning Experiments
K.1 BLEU Score for Flickr30k Experiments (Higher is better)
Figure 23: BLEU Score in lower supervision Flickr30k Image-Caption Prediction.
Figure 24: BLEU Score in higher supervision Flickr30k Image-Caption Prediction.
K.2 BLEU Score for COCO Experiments (Higher is better)
Figure 25: BLEU Score in lower supervision COCO Image-Caption Prediction.
Figure 26: BLEU Score in higher supervision COCO Image-Caption Prediction.
K.3 BLEU Score for Wiki Experiments (Higher is better)
Figure 27: BLEU Score in lower supervision Wiki Image-Caption Prediction.
Figure 28: BLEU Score in higher supervision Wiki Image-Caption Prediction.
Appendix L LLM Usage
@@ Appendix I Using Different Encoders and Clustering Methods
Appendix M More Information